OpenStudy (flygirlmandi):

which number is the solution of n divided by 3 equals -12 ? -36 -4 4 36 thanks :)

OpenStudy (alexandervonhumboldt2):

\[\frac{ n }{ 3 }=-12\] multply both sides by 3: \[\frac{ n }{ 3 }*3=-12*3\]

OpenStudy (alexandervonhumboldt2):

\[n=-12*3=???\]

OpenStudy (flygirlmandi):

so its a

alones (alones):

yes and negative multiplied by positive equals negative yes A

OpenStudy (flygirlmandi):

thanks guys :)